OpenStudy (anonymous):

Find the tangent line approximation for sqrt7+x near x=2.

OpenStudy (aum):

Let T(x) be the equation of the tangent line to the function f(x) at x = a. Slope of the tangent at x = a is f'(a) Therefore, \[T(x) = f(a) + f'(a)(x - a) \\ a = 2 \\ T(x) = f(2) + f'(2)(x - 2) \\ f(x) = \sqrt{7+x} \\ f(2) = \sqrt{7+2} = \sqrt{9} = 3\\ f'(x) = \frac {1}{2\sqrt{7+x}} \\ f'(2) = \frac {1}{2\sqrt{7+2}} = \frac 16\\ T(x) = 3 + \frac 16(x-2) \]

OpenStudy (anonymous):

Then multiply through and add 3?